OpenStudy (anonymous):

How do I find the real number solution? x³-3x²=0 I think I can factor by grouping to x^2(x-3) but then what?

OpenStudy (anonymous):

The zeros will be 0 and 3 because you will get x^2(x-3) = 0 x^2=0 x-3=0

OpenStudy (anonymous):

Oh ok. Thank you.

OpenStudy (anonymous):

Basically you once you have it grouped, you have to set it to zero to solve for the x value. When x-3 is solved, you get x=3 and x^2 =0 is already solved for x